OpenStudy (anonymous):

if 5x+1=21, then x = ?

OpenStudy (anonymous):

You need to get the variable on one side of the equation and the values on the other. How would you do this?

OpenStudy (arabpride):

start by combining like terms: 5x+1=21 -1 -1 5x = 20 divide by both sides: 5x = 20 /5 /5 x = 5 *^.^*

OpenStudy (anonymous):

x=4

OpenStudy (arabpride):

ooops /.^ my mistake yes that's correct x = 4